OpenStudy (anonymous):

PLEASE HELP!!!! Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11 i tried to solve for a but ended up with 1/-5 i dont think thats right....

OpenStudy (anonymous):



OpenStudy (anonymous):

a+b=4 2a+3b=11 a=4-b 2(4-b)+3b=11 (substituing a as found) a=4-b 8+b=11 a=4-b b=3 so, a=1 and b=3

OpenStudy (anonymous):

a=4-b 2(4-b)+3b=11 i get that but how did ypu go from that to a=4-b 8+b=11

OpenStudy (anonymous):

you need first to isolate a or b in one of the equations, and use this on the other. I have skipped one or two steps, but if you do correctly, the answer (a,b)=(1,3) will be found

OpenStudy (anonymous):

2(4-b) = 8-2b, write it and keep solving

OpenStudy (anonymous):

This problem is about solving what is called 'linear system'. You have some equations with letters and need to find what are their values. Further you will see that not all of this kind of systems can be solved (when there are more letters than equations) - but you can just ignore it by now tell me when you're done, it just a matter of practice to solve it better an faster

OpenStudy (anonymous):

ok question: in 2(4-b)+3b=11 would it equal 8-b+3b=11 or 8-2b+3b=11?

OpenStudy (anonymous):

8-2b+3b=11

OpenStudy (anonymous):

because the 2 multipies both 4 and -b in 2(4-b)

OpenStudy (anonymous):

this is what i got 8-b+3b=11 8+2b=11 2b=11-8 2b=3 2b/2=3/2

OpenStudy (anonymous):

re-do it haha it is 8-2b+3b=11 but what you've done is right if it was 8-b+3b=11

OpenStudy (anonymous):